### v1

# Modelling Sampling Distributions of Test Statistics with Autograd

Abstract
Simulation-based inference methods that feature correct conditional coverage of confidence sets based on observations that have been compressed to a scalar test statistic require accurate modelling of either the p-value function or the cumulative distribution function (cdf) of the test statistic. If the model of the cdf, which is typically a deep neural network, is a function of the test statistic then the derivative of the neural network with respect to the test statistic furnishes an approximation of the sampling distribution of the test statistic. We explore whether this approach to modelling conditional 1-dimensional sampling distributions is a viable alternative to the probability density-ratio method, also known as the likelihood-ratio trick. Relatively simple, yet effective, neural network models are used whose predictive uncertainty is quantified through a variety of methods.
1 Introduction
Automatic differentiation (see, for example, Ref.[1]) has revolutionized machine learning, permitting the routine application of gradient descent algorithms to fit to data models of essentially unlimited complexity. The same technology can be used to take the derivative of these models with respect to their inputs without the need to explicitly calculate the derivatives [2]. A potentially useful application of this capability is approximating the probability density function (pdf), $f(x\mid\boldsymbol{\theta})$, given an accurate neural network model of the associated conditional cumulative distribution function (cdf), $F(x\mid\boldsymbol{\theta})$, using the fact that
$f(x\mid\boldsymbol{\theta})$ $=\frac{\partial F(x\mid\boldsymbol{\theta})}{\partial x},$ (1)
where $\boldsymbol{\theta}$ are the parameters of the data-generation mechanism, which we distinguish from the parameters $\mathbf{w}$ of the neural network model. This paper explores this possibility in the context of simulation-based frequentist inference [3, 4, 5, 6, 7].
Equation (1) furnishes an approximation of the pdf $f(x\mid\boldsymbol{\theta})$ whether $x$ is a function of the underlying observations $D$ only or if $x=\lambda(D;\boldsymbol{\theta})$ is a test statistic that depends on $D$ as well as on the parameters $\boldsymbol{\theta}$. Moreover, computing the derivative of the cdf using autograd to obtain the pdf is exact; autograd does not use finite difference approximations.
Strictly speaking, Eq. (1) applies only if $x$ is from a continuous set. However, discrete distributions are frequently encountered in high-energy physics and other fields, and are often approximated by continuous distributions through suitable coarse-graining of $x$. Therefore, it is of interest to explore the degree to which we can approximate the cdf of discrete distributions with continuous ones and then apply Eq. (1) to obtain smooth approximations of the probability mass functions (pmf).
Given the pdf $f(x\mid\boldsymbol{\theta})$, where the data $D$ have been compressed to the 1-dimensional quantity $x$, further inferences can be performed by treating the pdf as a statistical model for $x$. These include constructing approximate confidence intervals by profiling  [8] $f(x\mid\boldsymbol{\theta})$ or constructing posterior densities for the parameters $f(\boldsymbol{\theta}\mid x)$ given a suitable prior. However, if $x=\lambda$ is a test statistic some care must be exercised because, in general, a variation of $\boldsymbol{\theta}$ will induce a change in $\lambda$, which should be taken into account. The key issue is whether a sufficiently accurate model of the cdf $F(x\mid\boldsymbol{\theta})$ can be constructed. This is the issue we explore in this paper where the focus is on test statistics that arise in simulation-based inference. We consider the classic ON/OFF problem of astronomy  [9] and high-energy physics as a benchmark example and apply the insights gained to the SIR problem [10, 11] in epidemiology.
This paper is structured as follows. In Sec. 2 we review related work in the areas of smiulation-based inference and uncertainty quantification in neural networks. In Sec. 3 we model the pdf from the cdf of a test statistic for problem where the likelihood is tractable: the prototypical signal/background problem in high energy physics, which in astronomy is known as the On/Off problem. In sec. 4 we model the pdf from the cdf of a test statistic for a problem where the likelihood is not tractable: the SIR model in epidemiology. In both examples, various techniques are applied to improve the accuracy and quantify the uncertainty. The paper ends with a discussion and conclusion in Sec. 5.
2 Related work
2.1 Likelihood Ratio Trick
It has been known since at least the early 1990s [12, 13, 14, 15] that probabilities can be modeled with neural networks using the density ratio method, also known as the likelihood ratio trick [3, 16, 17, 18, 19]. Suppose one has two data samples, $\{(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})\}_{A}$ and $\{(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})\}_{B}$. Sample $A$ comprises pairs $(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ in which the components are sampled sequentially: $\boldsymbol{\theta}_{i}\sim\pi$ followed by $\boldsymbol{x}_{i}\sim G(\boldsymbol{\theta}_{i})$, where $\pi$ is a known prior and $G$ is a simulator. Sample $B$ differs from $A$ in that $\boldsymbol{x}_{i}\sim g$, where $g(\boldsymbol{x}\mid\boldsymbol{\theta})$ is a known density that may, or may not, depend on $\boldsymbol{\theta}$. If sample $A$ is assigned target $t=1$ and $B$ is assigned $t=0$, then a sufficiently flexible function — fitted by minimizing the cross-entropy loss averaged over a large balanced training sample — yields an approximation to the function
$D(\boldsymbol{x},\boldsymbol{\theta})$ $=\frac{f(\boldsymbol{x},\boldsymbol{\theta})}{f(\boldsymbol{x},\boldsymbol{\theta})+g(\boldsymbol{x}\mid\boldsymbol{\theta})\pi(\boldsymbol{\theta})},$ (2)
where $f(\boldsymbol{x},\boldsymbol{\theta})$ is the (generally unknown) joint density associated with sample $A$. A rearrangement of Eq.(2) leads to the result
$f(\boldsymbol{x}\mid\boldsymbol{\theta})$ $=\frac{f(\boldsymbol{x},\boldsymbol{\theta})}{\pi(\boldsymbol{\theta})},$
$=g(\boldsymbol{x}\mid\boldsymbol{\theta})\left(\frac{D}{1-D}\right).$ (3)
This method and related methods are available in the Madminer package [5].
2.2 LF2I and ALFFI
Our motivation for starting with the cdf (or p-value) is that constructing a model of it is the key step in the likelihood-free frequentist (LF2I) approach  [6] and in an extension of it (ALFFI)  [7] in which the test statistic $x=\lambda(D;\boldsymbol{\theta})$ is included as an input to the neural network-based model of the cdf. If the test statistic $\lambda$ is chosen so that large values of $\lambda$ disfavor the hypothesis $H:\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$ it follows that a confidence set, $R(D)$, can be constructed at confidence level $\tau=1-\alpha$ where $\alpha$ is the miscoverage rate. By definition the set $R(D)$ is all values of $\boldsymbol{\theta}_{0}$ for which $F(\lambda\mid\boldsymbol{\theta}_{0})\leq\tau$ given data $D$ [6, 7]. The fact that $\lambda$ is an input to the model in the ALFFI algorithm presents an opportunity: taking the derivative of the approximate cdf $\hat{F}$ with respect to $\lambda$ provides an approximation $\hat{f}(\lambda\mid\boldsymbol{\theta})$ of the sampling distribution of the test statistic. In this paper we explore the accuracy with which both the cdf and pdf can be approximated.
2.3 Conformal Inference
Suppose we have $n$ training samples $\left(\mathbf{x}_{i},y_{i}\right)\in\mathbb{R}^{d}\times\mathbb{R},i\in[1,n]$, where $x_{i}$ is a $d$-dimensional feature vector and $y_{i}$ is the response variable (target). Let $f(x)$ denote the regression function (such as a neural network), which is often fitted by minimizing the average quadratic loss between the target and the function, in which case $\hat{f}(x)\approx\mathbb{E}[y\mid x]$. We are interested in predicting a new response $y_{n+1}$ from a new feature vector $x_{n+1}$. Given a miscoverage rate $\alpha\in[0,1]$ we wish to build a confidence set $\mathcal{C}_{\alpha}$ with the property that
$\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha=\tau.$ (4)
The confidence set is a measure of uncertainty in the prediction of the function $f(x)$.
[…]
$\mathcal{C}_{\alpha}=[\hat{f}(x_{n+1})-\hat{q}_{1-\alpha},\ \hat{f}(x_{n+1})+\hat{q}_{1-\alpha}].$ (5)
More details on conformal prediction are provided in [25, 26]. We use conformal inference to construct confidence sets for the cdf and pdf of $\lambda$ as a way to quantify their accuracy and for potentially correcting the cdf to arrive at a more accurate cdf model.
2.4 Multistage Modelling of Neural Networks
Recently, multi-stage neural networks (MSNN) [27] (see also [28]) has been proposed as a strategy to approximate the target function of neural networks with remarkable accuracy, with the prediction errors approaching machine precision $\mathcal{O}(10^{-16})$ for double floating point numbers. The method constructs a sequence of neural networks, each fitted to the residuals from the previous stage. We follow the notation in [27].
Let $x$ be the input features and $u_{g}(x)$ be the target function. The data $(x,u_{g}(x))$ are used to train a neural network $u_{0}(x)$ to regress $u_{g}(x)$. The error, or residual, between the neural network and the target function $e_{1}(x)=u_{g}(x)-u_{0}(x)$ is calculated. The residual data $(x,e_{1}(x)/\epsilon_{1})$ are used to train a second neural network $u_{1}(x)$ to regress $e_{1}(x)/\epsilon_{1}$, where $\epsilon_{1}$ is the root-mean square,
$\epsilon_{1}=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left[e_{1}\left(x\right)\right]^{2}}=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left[u_{g}-u_{0}\left(x\right)\right]^{2}}.$ (6)
The target function is rescaled so that its range is of order one. The corrected regression function is given by
$u_{c}^{(1)}(x)=u_{0}(x)+\epsilon_{1}u_{1}(x).$ (7)
By fitting a second neural network on the residuals and adding it to the original network one improves the precision of the regression function. The algorithm can be continued to train further neural networks on $\left(x,e_{n}\left(x\right)/\epsilon_{n}\right)$ to reach higher accuracy and all $(n+1)$ neural networks are combined as follows,
$u_{c}^{(n)}(x)=\sum_{j=0}^{n}\epsilon_{j}u_{j}(x),$ (8)
where $\epsilon_{i}$ refers to the root mean square for the $i$-th neural network with $\epsilon_{0}=1$. We explore whether the cdf to be modeled is smooth enough for this method to work (see Sec. 3.3).
2.5 Bootstrap NNs
The bootstrap [29] is a statistical method that treats an observed dataset as if it were a population. This makes it possible, for example, to approximate the sampling distributions of statistics. Given an estimator $\hat{\theta}$ of $\theta$ the bootstrap quantifies the uncertainty in $\hat{\theta}$ by repeatedly sampling the original dataset with replacement to create new datasets. The uncertainty in $\hat{\theta}$ can be quantified with a measure of the variability of predictions between different bootstrap datasets [30, 31].
In the context of supervised learning, the samples are given in pairs $z_{i}=(\mathbf{x}_{i},y_{i})$ from a joint distribution $p_{\theta}(\mathbf{x},y)$. Suppose we are given a set of features and targets composed of $n$ examples $\mathbf{z}=\{\left(\mathbf{x}_{i},y_{i}\right)\}_{i=1}^{n}$. One draws a bootstrap dataset from the original training dataset of the same size as the original dataset. This is done $K$ times and the same model is fitted to each of the $K$ bootstrap datasets using the same training protocol. This yields $K$ neural networks and, therefore, $K$ outputs for a given input. A measure of the spread of the outputs quantifies the uncertainty in the model output.
3 Example 1: signal/background or ON/OFF model
For our first example, we choose a problem that is ubiquitous in high-energy physics and astronomy: the signal/background problem, also known as the ON/OFF problem in astronomy [9] and we consider its simplest realization. An observation is made which consists of counting $N$ events (these are particle collisions in particle physics or photon counts in astronomy). A second independent observation is made where no signal is present by design, yielding $M$ counts. Following [7] the statistical model is taken be a product of two Poisson probability mass functions (pmf) for the observed data $D=\{N,M\}$,
$\mathbb{P}(N,M\mid\mu,\nu)=\mathcal{L}(D;\mu,\nu)$ $=\frac{(\mu+\nu)^{N}\exp(-(\mu+\nu))}{N!}\,\frac{\nu^{M}\exp(-\nu)}{M!},$ (9)
where $\mu$ and $\nu$ are the mean signal and background counts, respectively, and we use the same likelihood ratio test statistic as in [7]
$\lambda(D;\mu,\nu)=-2\log\left[\frac{\mathcal{L}(D;\mu,\nu)}{\mathcal{L}(D;\hat{\mu},\hat{\nu})}\right],$ (10)
where the maximum likelihood estimate of the signal is $\hat{\mu}=N-M$, which can be positive or negative. Negative signal estimates are avoided by using the “non-maximum likelihood estimate”
$\hat{\mu}=\left\{\begin{array}[]{ll}N-M&\text{ if }\quad N>M\\ 0&\quad\textrm{ otherwise.}\end{array}\right.$ (11)
and
$\hat{\nu}=\left\{\begin{array}[]{ll}M&\text{ if }\quad\hat{\mu}=N-M\\ (M+N)/2&\quad\textrm{ otherwise.}\end{array}\right.$ (12)
Since $\mu$ is the parameter of interest and $\nu$ is a nuisance parameter one is generally interested in confidence intervals for $\mu$ regardless of the true value of $\nu$. However, the sampling distribution of $\lambda$, $f(\lambda\mid\mu,\nu)$, in general depends on both the parameter of interest and the nuisance parameter. This remains true even if we were to replace Eq. 10 with the profile likelihood ratio in which the nuisance parameter $\nu$ is replaced by its conditional estimate. This is because of the exclusion of negative estimates of the signal. Therefore, we still must contend with the nuisance parameter.
There are at least two plausible ways one might proceed given an approximation to the sampling distribution, $f(\lambda\mid\mu,\nu)$: either replace the nuisance parameter by a plug-in estimate or replace it through a procedure analogous to profiling but applied to $f(\lambda\mid\mu,\nu)$. However, while profiling one has to account for the fact that $\lambda(D;\mu,\nu)$ depends on the nuisance parameter. The viability of using a profiled sampling distribution for creating approximate confidence intervals for $\mu$ alone remains to be explored.
3.1 Modelling the CDF with ALFFI
ALFFI is an algorithm for approximating the conditional cdf, $F(x\mid\theta)$, of a scalar random variable $x$ where $\theta$ denotes the parameters of the statistical model. If $x=\lambda(D;\theta)$ is a test statistic then the cdf can be used to construct confidence sets for all parameters simultaneously. Here we explore the accuracy of the ALFFI algorithm for modelling the cdf with a view to deriving the sampling distribution, $f(\lambda\mid\mu,\nu)$ of the ON/OFF test statistic by differentiating the cdf with respect to $\lambda$.
Again following [7] we start by sampling $\mu$ and $\nu$ from uniform priors and sample $n_{i}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $m_{i}\sim\text{Poiss}(\nu_{i})$. At each parameter point, we calculate $\lambda_{i}=\lambda(n_{i},m_{i}\mid\mu_{i},\nu_{i})$ according to Eq. 10 and the procedure is repeated to sample $N_{i},M_{i}$ and $\lambda_{D}=\lambda(N_{i},M_{i}\mid\mu_{i},\nu_{i})$. Finally, the indicator $Z$, which is unity if $\lambda_{i}\leq\lambda_{D}$ and zero otherwise, is computed. This results in a training set of size $B$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},Z_{i})\}_{i=1}^{B}$. The observation in ALFFI and LF2I is that critical value functions, such as the cdf $F(\lambda\mid\mu,\nu)=\mathbb{P}(\lambda\leq\lambda_{D}\mid\mu,\nu)$, are the expectation value $\mathbb{E}(Z\mid\lambda_{D},\mu,\nu)$ of the discrete random variable $Z$, of which a smooth approximation can be created with a deep neural network $f(\mathbf{x}_{i},\mathbf{w})$. Critically, this network must be trained to minimize the mean square error (MSE) loss [^1] In LF2I and ALFFI, the predicted value is $\mathbb{E}[Z\mid x]=\mathbb{P}(Z=1\mid x)=\frac{\mathbb{P}(x\mid Z=1)\mathbb{P}(Z=1)}{\mathbb{P}(x\mid Z=1)+\mathbb{P}(x\mid Z=0)}$. In other words, it behaves like a classifier, and because of that the cross entropy loss can also be used.,
$L(\mathbf{w})=\frac{1}{B}\sum_{i=1}^{B}\left(y_{i}-f(\mathbf{x}_{i},\mathbf{w})\right)^{2},$ (13)
where $\mathbf{w}$ are the parameters (weights) of the neural network and $\mathbf{x}_{i}=\{\mu_{i},\nu_{i},\lambda_{i}\}$ is a batch of training data, with targets $y_{i}=Z_{i}$. Details of this algorithm are provided in Appendix A.
The cdf, using the ALFFI algorithm, was approximated using a fully connected neural network with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$, 12 hidden layers with 12 nodes each, and a single output. A sigmoid was used in the output layer to constrain the output to the unit interval. The activation function at each hidden node is a SiLU [32]. The SiLU nonlinearity was used in all the models in this study, since the SiLU allows the model to be smooth and differentiable multiple times. The network was trained with the NAdam optimizer [33] with a fixed learning rate of $6\times 10^{-4}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $512$ for $10^{5}$ iterations, that is, for 5 epochs. Longer training runs were also performed, but did not yield improvements in the results.
Figure 1: ON/OFF cdfs modelled with ALFFI at different $(\mu,\nu)$ points.
Figure 2: ON/off pdfs obtained by differentiating cdfs modelled with ALFFI at different $(\mu,\nu)$ points.
Although the ALFFI algorithm yields valid confidence sets for the ON/OFF problem, it does not yield a sufficiently accurate approximation of the cdf, and hence of the pdf. This is shown in Fig. 1, where the empirical cdf [^2] Suppose that we have a sample of observations $\lambda_{1},...,\lambda_{n}$. The empirical cdf is defined by $F(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(\lambda_{i}\leq\lambda\right)$, $F$, is displayed together with the approximated cdf $\hat{F}_{\text{ALFFI}}(\lambda\mid\mu,\nu)$. Taking the derivative of the approximated cdf yields the ALFFI approximation of the pdf, $\hat{f}_{\text{ALFFI}}(\lambda)=\frac{\partial\hat{F}_{\text{ALFFI}}}{\partial\lambda}$, displayed in Fig. 2. We observe that $\hat{f}_{\text{ALFFI}}(\lambda)$ displays sharp fluctuations and turning points, which was also confirmed by taking a numerical derivative of $\hat{F}_{\text{ALFFI}}$. This is because the actual slope of the cdf changes dramatically from point to nearby point, especially at low values of $\lambda$. It can also be seen that the histogrammed pmf of $\lambda$ at a particular $(\mu,\nu)$ point is a high frequency function, or one that is noisy and features very sharp peaks. The fluctuations seen in the pmf also depend on the choice of binning: finer binning shows larger fluctuations because this is a discrete distribution. This high-frequency nature of the pmf complicates our attempt to create a smooth modelling of the pmf with ALFFI.
3.2 Directly modelling the empirical cdf
Figure 3: ON/OFF cdfs from modelling the empirical cdf with the latter as the targets at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band attained from conformal inference.
Figure 4: ON/OFF pdfs obtained by differentiating the models of the empirical cdf at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band attained from conformal inference..
Trying to model an intrinsically discrete distribution with a smooth approximation may be beyond an algorithm such as ALFFI and one might argue it is in any case a hopeless task. Nevertheless it is useful to check if an alternative, more direct, approach works better: directly modelling the empirical cdf. At each $\{\mu_{i},\nu_{i}\}$ point, we generate $K=100$ experiments: $\{N_{i}^{j}\}_{j=1}^{K}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $\{M_{i}^{j}\}_{j=1}^{K}\sim\text{Poiss}(\nu_{i})$, for which we generate $K$ instances of the test statistic $\{\lambda_{i}^{j}\}_{j=1}^{K}$ according to Eq. 10. These data are used to calculate the empirical cdf $F_{i}(\lambda_{i})$ at every parameter point. The data were then flattened to result in a training set of size $B=10^{7}$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},F_{i})\}_{i=1}^{B}$.
The cdf was again approximated with a neural network with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$ but this time with the targets set to values of the empirical cdf $F$. The architecture consisted of 6 hidden layers with 12 nodes each and a single output. The activation function at each hidden node is a SiLU. The network was trained to minimize the mean square error with the NAdam optimizer with a fixed learning rate of $3\times 10^{-5}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $60$ for $10^{6}$ iterations, that is, for 6 epochs.
A much better approximation of the cdf is obtained, as shown in Fig. 3 and more importantly in the associated pdf. Fig. 4 shows that the jagged fluctuations in the pdf disappear and we obtain a smooth approximation of the discrete distribution. Modelling the empirical cdf presents yet another advantage: the possibility of using the empirical cdf as the true response variable in the conformal inference algorithm (see Sec. 2.3) and thereby providing a quantification of the uncertainty in the modelling of the cdf and pdf. We use conformal inference to calculate a $68\%$ confidence interval in the cdf space for each value of the test statistic. For the associated pdf, a coarse-graining of $\lambda$ is performed by histogramming $\lambda$ and the coarse-grained $f(\lambda)$ was used as the true response variable (that is, the true pdf). The excellent agreement between the predicted pdf and the histogrammed pdf is shown in Fig. 4 along with the $68\%$ conformal confidence interval at each value of $\lambda$.
3.3 Multistage Modelling of the CDF
Another approach to improve the accuracy of regression functions is described in [27], which proposes a multi-stage neural network (MSNN) strategy in which the network training is divided into multiple stages, where each stage fits a separate neural network to the residuals from the previous stage (see Sec. 2.4).
Quantile residuals ($\Delta C$): Let $p=\{F_{1},...,F_{k}\}$ be a set of evenly spaced probabilities and let $\hat{F}=\{\hat{F}_{1},...,\hat{F}_{n}\}$ be a set of estimated cdf values at a given $(\mu,\nu)$ parameter point with $n>>k$. Define $q$ as the quantiles of $\hat{F}$ associated with the probabilities $p$. If we can model the correction function $p=C(q)$ that maps the estimated cdfs $q$ to the exact cdf $p$ with sufficient accuracy then, in principle, the correction function can be used to improve the estimated cdf. Writing $p=q+\Delta C(q)$, we define the quantile residual function as $\Delta C=q-p$ and try to model it as a function of $q$ following the multistage method.
In order to follow the multistage method for constructing a model of the cdf one needs an accurate model of the residuals. We anticipate that this will be challenging because the distribution of $\lambda$ is discrete and their values are not evenly spaced. Indeed, this is borne out in Fig. 9 which shows that $\Delta C$ is a noisy high frequency function. Furthermore, as can be seen in Fig. 9, the functions do not seem to vary smoothly with the parameters $(\mu,\nu)$ (small changes in the parameter values yield large changes in the residual function).
Nevertheless, we wish to see whether a neural network can produce a smooth interpolation of this noisy function by implementing the suggestions described in [27]: 1. Use a sine nonlinearity in the first layer and tanh nonlinearity in subsequent layers. 2. Use Glorot weight initialization [34] with a constant factor $\kappa=60$ multiplying the weights in the first layer and no such factor in subsequent layers. It was observed that the $\kappa$ weight factor helps to stabilize the training and the sine nonlinearity helps capture the high frequency. 3. Normalize the target function by the RMS of the outputs of the target function $\Delta C\rightarrow\Delta C/RMS(\Delta C)$, as in Eq. 6. Further experiments were performed such as adding or omitting dropout, batch normalization, weight regularization, and experimenting with different nonlinear functions such as ReLU, SiLU, tanh, sine, and groupsort [35]. None of these experiments were able to capture the high frequency behavior of the residual target function when trained on the residuals of all parameters in the training data simultaneously. However, we were able to use a simple neural network to model the residual function when modeled point-by-point in the parameter space.
The frequency principle [36] states that neural networks first fit low frequencies and later fit high frequencies during the training process. However, prior studies [37, 38] have demonstrated that in practice standard neural networks struggle to fit high frequency functions. We confirm these results and observe that the neural networks we used struggle to fit the residual functions even when trained for very long training runs. LSTMs, [39] on the other, hand are known to outperform standard neural networks in fitting high frequency and sequential data, where the order of the data is significant [40]. They have been successfully applied in many areas in sequence learning and time series forecasting such as financial forecasting [41], handwriting recognition [42] and speech recognition [43]. In our study, LSTMs achieve reasonably good predictions of the residuals when trained over all $(\mu,\nu)$ values. Although our LSTM models outperform our multi-layer perceptron (MLP) models in this task the results are unsuitable for use in the MSNN approach due to inadequate precision in the resulting predictions.
Cdf residuals ($\Delta r$): An alternative to modelling the quantile residuals is to model the cdf residuals as a function of $\lambda$, which is closer to what is done in the MSNN approach. At a particular $(\mu,\nu)$ point, the cdf residual is simply the difference between the predicted and the empirical cdf: $\Delta r=\hat{F}-F$. Figure 10 shows that the cdf residuals as a function of $\lambda$ are a slightly smoother function of the parameter points than is $\Delta C$. Nevertheless, the MLP and LSTM models still struggle to capture the high-frequency nature of this residual function. A coarse-graining of $\lambda$ was also performed to reduced the high-frequency of the residual function but to no avail.
4 Example 2: SIR Model
The lessons learned in the ON/OFF example are applied to modelling the cdf in a different context: epidemiology. We use the SIR model [10, 11] which treats an epidemic as a system in which transitions occur between three states or compartments: susceptible (S), infected (I), and recovered (R). This model has both a stochastic realization as well as a description in terms of ordinary differential equations
$\frac{dS}{dt}$ $=-\beta\,SI,$
$\frac{dI}{dt}$ $=-\alpha I+\beta\,SI,$
$\frac{R}{dt}$ $=\alpha I,$ (14)
whose solutions approximate the mean number of susceptible $(S(t))$, infected ($I(t)$) and recovered ($R(t)$) individuals as a function of time, $t$. The model depends on two parameters $\alpha$, the rate of recovery, and $\beta$, the rate of transmission per infected individual. We fit the model to the data $D=\{x_{1},...,x_{n}\}$ by minimizing the following test statistic [7]
$\lambda(D;\theta)$ $=\frac{1}{50}\sqrt{\frac{1}{N}\sum_{n=1}^{N}\frac{[x_{n}-I_{n}]^{2}}{I_{n}}},$ (15)
where $\theta=\{\alpha,\beta\}$, $x_{n}$ is the observed number of infected individuals at the observation time $t_{n}$ and $I_{n}=I(t_{n},\theta)$ is the predicted mean infected count obtained by solving the system of ordinary differential equations, Eq. 14. This example is chosen because it exemplifies the utility of the methods discussed in this paper for inference with intractable statistical models. All of our major findings from the ON/OFF example apply to this example, but here we wish to improve the modelling of the cdf and employ more techniques for uncertainty quantification.
4.1 Directly modelling the empirical cdf
One of the lessons from the ON/OFF problem is that for discrete distributions the ALFFI algorithm does not yield a sufficiently accurate smooth model of the cdf, which was confirmed in the SIR example. We therefore directly model the empirical cdf as a function of the model parameters $(\alpha,\beta)$ and the test statistic. The training dataset comprises $250$ uniformly sampled $(\alpha,\beta)$ parameter points, where at each $(\alpha_{i},\beta_{i})$ point, $K=400$ epidemics were simulated yielding a set of test statistics $\{\lambda_{i}^{j}\}_{j=1}^{K}$ and a set of empirical cdfs $\{F_{i}^{j}\}_{j=1}^{K}$. The data were then flattened to result in a training set of size $B=10^{7}$, $\mathcal{T}=\{(\alpha_{i},\beta_{i},\lambda_{i},F_{i})\}_{i=1}^{B}$.
The MSE loss is known to be sensitive to outliers in the data. Therefore, we chose to use the more robust Huber loss [44], which is an MSE loss for relatively small errors and an absolute loss for larger errors. The Huber loss is given by
$L_{\delta}(\boldsymbol{\omega})=\left\{\begin{array}[]{ll}\frac{1}{2}(y-f(\mathbf{x},\boldsymbol{\omega}))^{2}&\text{for }|y-f(\mathbf{x},\boldsymbol{\omega})|\leq\delta\\ \delta\left(|y-f(\mathbf{x},\boldsymbol{\omega})|-\frac{1}{2}\delta\right),&\text{otherwise.}\end{array}\right.$ (16)
where $\delta$ is a tunable parameter. The Huber loss is both differentiable everywhere and robust to outliers. We employ models trained with MSE and models trained with the Huber loss and observe that for relatively short training runs ($\approx 10^{5}$ iterations) the Huber loss with $\delta=0.7$ significantly outperforms the MSE loss when all other hyperparameters of the model are held fixed. For larger training runs ($>10^{6}$ iterations), however, using the Huber loss and MSE loss yields approximately the same results. In order to reduce the number training iterations, we use the model with Huber loss for all of our subsequent models.
All of the models for this example approximate the cdf using a fully-connected neural network with 3 input features $\mathbf{x}=\{\alpha,\beta,\lambda\}$, 5 hidden layers with 10 nodes each, and a single output using a sigmoid. The activation function at each hidden node is a SiLU. The network was trained with the NAdam optimizer with a fixed learning rate of $3\times 10^{-4}$. The training set, which comprises $10^{7}$ examples, was used in batches of size $60$ for $3\times 10^{6}$ iterations, that is, for 18 epochs. As the training proceeds the network with the lowest validation loss is saved. Figures 5 and 6 show the resulting estimated cdfs and pdfs with the 68% conformal confidence intervals at different values of $\lambda$ and for different $(\alpha,\beta)$ points. A coarse-graining of $\lambda$ was performed as a way to estimate the true pdf in order to apply the conformal inference algorithm.
Figure 5: Cdfs for the SIR model obtained from modelling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band.
Figure 6: Pdfs for the SIR model obtained from modelling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band. The spikes around $\lambda\approx 0.15$ arise from simulated epidemics that die off within a day or so. In a considerably larger sample of simulated epidemics the spikes would not be visible.
4.2 Uncertainty quantification using Bayesian neural networks
Suppose we have a training dataset $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ (where $\mathbf{x}$ and $y$ are assumed to be drawn from a joint distribution $p(\mathbf{x},y)$). Our goal is to predict the target variable $y$ given a new input value $\mathbf{x}$. From a Bayesian perspective, the goal is to approximate the posterior predictive distribution $p(y\mid\mathbf{x},\mathcal{D})$ given new input $\mathbf{x}$ and training data $\mathcal{D}$,
$p(y\mid\mathbf{x},\mathcal{D})=\int p(y\mid\mathbf{x},\mathbf{w})\,p(\mathbf{w}\mid\mathcal{D})\,\mathrm{d}\mathbf{w},$ (17)
which entails a marginalization over the network parameters. By Bayes’ theorem, the posterior density over the neural network parameter space is given by
$p(\mathbf{w}\mid\mathcal{D})=\frac{p(\mathcal{D}\mid\mathbf{w})p(\mathbf{w})}{p(\mathcal{D})},$ (18)
which requires the model evidence,
[…]
Unlike the best-fit approach to neural networks, a Bayesian neural network (BNN) is the posterior density $p(\mathbf{w}\mid\mathcal{D})$ over the network parameter space.
The severe bottleneck with BNNs is computing the high-dimensional integrals Eq. 17 and Eq. 19. In practice, these integrals must be approximated using various sampling schemes including Hamiltonian Monte Carlo [45] or variational inference [46]. Furthermore, a prior density $p(\mathbf{w})$ over the neural network parameter space must be specified. Given a collection of $K$ neural networks with parameters sampled from the posterior density, the predictive distribution can be approximated as follows
$p(y\mid\mathbf{x},\mathcal{D})$ $=\frac{1}{K}\sum_{k=1}^{K}p(y|\mathbf{x},\mathbf{w}_{k}),$ (20)
where, typically, one models the density $p(y|\mathbf{x},\mathbf{w}_{k})$ as
$p(y|\mathbf{x},\mathbf{w}_{k})\approx\mathcal{N}(y;f(\mathbf{x},\mathbf{w}),\sigma),$ (21)
with $f(\mathbf{x},\mathbf{w})$ the neural network.
Figure 7: Fluctuating NN parameters by sampling from a prior modeled as a multivariate diagonal normal distribution with a single standard deviation $\sigma$. The spread of results is seen to be particularly sensitive to the choice of $\sigma$. We speculate that this could be due to the lack of adaptation of $\sigma$ to the different directions in the neural network parameter space.
In practice, given the computational burden of sampling from the true posterior density, we adopt a simpler approach. Let $\mathbf{w^{*}}\in\mathbb{R}^{k}$ be the best-fit neural network parameters obtained by our training protocol. We fluctuate the neural network parameters by sampling $\mathbf{w}_{1},...,\mathbf{w}_{n}\sim\mathcal{N}(\mathbf{w^{*}},\sigma^{2}\mathbf{I}_{k})$ for different choices of $\sigma$. With $n=100$, the resulting ensemble of models evaluated at the sampled parameters $\hat{f}(\mathbf{w}_{1}),...,\hat{f}(\mathbf{w}_{n})$ are shown in Fig. 7.
4.3 Bootstrap NN
We follow the bootstrap idea (see Sec. 2.5) to draw datasets with replacement from the training data, such that each bootstrap sample has the same size as the original training set. This is done $K=200$ times, resulting in $K=200$ bootstrap training datasets. The same model with the architecture described in Sec. 4.1 is re-trained for $10^{5}$ iterations, with each model fitted to a different bootstrap sample. The spread of the resulting models over the $K=200$ neural networks is displayed in Fig. 8. The envelope of the spread is observed to be much smaller than for the previous methods. This could be due to insufficient training of the models. Other possible reasons for this are outlined below.
Figure 8: Responses of the 200 neural networks each trained on a different bootstrap training sample. See Sec. 4.3 for details.
Although the bootstrap has been hailed as a satisfactory method for uncertainty quantification due to its theoretical guarantees [47] it has limitations. For example, it is known to fail in situations where the data have a non-trivial dependency structure or when the functionals of interest are not smooth [48].
Another question that was recently explored is the efficacy of bootstrap in the high-dimensional regime $n/d<1$ relevant to machine learning, where $d$ is the dimension of the machine learning model parameter space and $n$ the training sample size. In [49] it is shown that even for the simple case of inference in the linear model with unregularized least squares, bootstrap techniques perform poorly in only moderately high dimensions even with $n>d$. It was also shown in [50] that resampling methods such as the bootstrap yield reliable error estimates only in the very low-dimensional regime $n>>d$, and is fraught with problems in high-dimensional regimes. Therefore, perhaps one shouldn’t be surprised by the results described above.
5 Discussion and conclusion
In this paper we try to answer the question whether one can accurately model the sampling distribution of a test statistic starting from a neural network model of the cdf.  [^3] Code to reproduce all of our results is available at https://github.com/hbprosper/cdf2pdf.. Our motivation for modelling the cdf $F(\lambda\mid\boldsymbol{\theta})$ as opposed to an ab initio modelling of $f(\lambda\mid\boldsymbol{\theta})$ is that in general a better precision is reached approximating integrals of functions rather than the functions themselves. Furthermore, since the step of taking the derivative with autograd is exact, the accuracy of the approximated cdf directly translates to the accuracy of the desired pdf.
It was found that a recently introduced method (ALFFI) that approximates the cdf as the mean of a certain discrete random variable appears not to be accurate enough in modelling the cdf for the purpose of deriving its associated pdf. We caution however that this may be the inevitable consequence of trying to model the cdf of an intrinsically discrete distribution with a smooth approximation.
We employ relatively simple yet effective neural network models such that the amount of fine-tuning that is involved in going to more complex data settings is minimal. Several neural network uncertainty quantification techniques were reviewed and implemented and we studied possible corrections of the cdf. Of the uncertainty quantification methods considered only one is calibrated by construction, namely, conformal inference. We recommend that it be used as a simple benchmark with respect to which other uncertainty quantification methods can be compared and calibrated.
As can be seen in Eq. 4 conformal inference results in confidence sets that guarantee marginal coverage, $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha$, that is, the coverage marginalized (integrated) over the population of which a dataset is a sample. That property is not at all controversial as coverage is a property of the population from which the dataset is presumed to have been sampled. But if each element $x$ of the dataset contains a subset that are parameters $\theta$ sampled from a prior $\pi_{\theta}$ then the marginalization is also with respect to this prior. When a prior is involved this form of coverage is weaker than the more sought after conditional coverage guarantee $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\mid\theta\right\}\geq 1-\alpha$. It was shown in [23] that such distribution-free conditional coverage is impossible to achieve with a finite sample, which has inspired studies that try to bridge the gap between marginal and conditional coverage [51]. In our case, $x=\{\theta,\lambda\}$ and we generate calibration sets at each $\theta$ point, thereby making our conformal confidence sets adaptive, that is, change with $\theta$. However, they are still not adaptive with respect to $\lambda$; the conformal confidence intervals are the same width at each $\lambda$ value.
Another issue is that our conformal prediction sets are adaptive only because we have calibration data at the desired parameter points. It would be helpful to have a smooth interpolation of the width of the conformal confidence interval so that an interval can be computed at points where no calibration data exist.
[…]
Acknowledgements This work is supported in part by the US Department of Energy under Award No. DE-SC0010102 and in part by the University Research Association (URA) Visiting Scholars Program Award No. 23-S-01. We are deeply thankful to Pushpalatha Bhat and Stephen Mrenna for their insightful feedback on this project and their patience. Special thanks to the Fermilab LHC Physics Center (LPC) for providing access to resources and facilities that were instrumental to the completion of this work.
Appendix A ALFFI algorithm for modelling CDF for ON/OFF Example
ALFFI algorithm for modelling CDF in the On/Off problem, inspired by [7].
Algorithm 1 Estimate the CDF $\mathbb{C}(\lambda\mid\mu,\nu)$, given the observed data $\{N,M\}$ and the observed test statistic $\lambda$.
1:estimated CDF ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$ for all $\theta=\theta_{0}\in\Theta$
2:set $\mathcal{T^{\prime}}\leftarrow\emptyset$
3:for $i$ in $\{1,...,B^{\prime}\}$ do
4:     Draw parameter $\mu_{i}\sim\pi_{\mu}=\text{Unif}(0,20)$
5:     Draw parameter $\nu_{i}\sim\pi_{\nu}=\text{Unif}(0,20)$
6:     Draw $n_{i}\sim\text{Poiss}(\mu_{i}+\nu_{i})$
7:     Draw $m_{i}\sim\text{Poiss}(\nu_{i})$
8:     Draw parameter $\mu_{i}^{\prime}\sim\pi_{\mu^{\prime}}=\text{Unif}(0,20)$
9:     Draw parameter $\nu_{i}^{\prime}\sim\pi_{\nu^{\prime}}=\text{Unif}(0,20)$
10:     Draw $N_{i}\sim\text{Poiss}(\mu_{i}^{\prime}+\nu_{i}^{\prime})$
11:     Draw $M_{i}\sim\text{Poiss}(\nu_{i}^{\prime})$
12:     Compute test statistic under the null $\lambda_{i}\leftarrow\lambda(n_{i},m_{i}\mid\mu_{i},\nu_{i})$
13:     Compute indicator $Z_{i}\leftarrow\mathbbm{1}(\lambda_{i}\leq\lambda(N_{i},M_{i}\mid\mu_{i},\nu_{i}))$
14:     $\mathcal{T^{\prime}}\leftarrow\mathcal{T^{\prime}}\cup\{(\mu_{i},\nu_{i},\lambda_{i},Z_{i})\}$
15:end for
16:Use $\mathcal{T^{\prime}}$ to learn the parameterized function ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu):=\mathbb{E}[Z\mid\mu,\nu]$ via regression of $Z_{i}$ on $\{\mu_{i},\nu_{i},\lambda_{i}\}$ using mean square error as the loss function
17:return ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu)$, which estimates $\mathbb{P}(\lambda_{i}\leq\lambda(N,M\mid\mu,\nu))$
Appendix B Residuals
Figure 9: $\Delta C=\hat{q}-p$ residuals as a function of $\hat{q}$ at different $(\mu,\nu)$ points.
Figure 10: $\Delta r=\hat{F}-F$ a a function of $\lambda$ at different $(\mu,\nu)$ points.